A tobramycin level is determined at 3:00pm, 30 minutes after the end of a 30 minute IV infusion (started at 2:00 pm). Another level is drawn at 6:45pm. The two levels are 6.2μg/cc and 4.2μg/cc, respectively,

a). What is the t1/2 b). What is the patient's volume of distribution (in L/kg) [if dose =80mg/0.5hr and patient weight =70kg]
Answer #1

A= Ao e^-kt

4.2 = 6.2 e^(-k* 3.75 hour)

k= 0.22/ 3.75 =0.06

So t1/2 = 0.693 / k = 0.693/0.06 = 11.55 hours= 12 hours

Dose = 80 mg= 80000 micrograms

Concentration = 6.2 microgram/cc

So Vd = 80000/6.2 = 12903 cc

1 cc = 1 ml

So Vd = 12.9 L

=12.9/ 70 = 0.18 L/kg